# Thermal description of braneworld effective theories

### v1

@@ Abstract
A natural question to ask is- what is the low-energy effective theory on the branes? Since both cosmological and astrophysical solutions occur at energy scales much lower than that of the Planck scale, therefore in the effective theory approach the brane curvature radius $L$ is much large compared to bulk curvature $l$. As a result, perturbation theory can be used with a dimensionless perturbation parameter $\epsilon$ such that $\epsilon=(\frac{l}{L})^{2}<<1$. This method, called the Kanno-Soda gradient approximation scheme [4], is a metric-based iterative method in which the bulk metric and extrinsic curvature are expanded with increasing order of $\epsilon$ in perturbation theory. The effective Einstein equations on a brane are determined with the solutions of these quantities and the junction conditions. When this method is applied to the two-brane Randall-Sundrum (RS) model [3], the RS fine-tuning condition (see Eq.11) is reproduced at the zeroth order when the radion is constant and the two $3$-branes are characterised by opposite brane tensions. The effective Einstein equations are then obtained at the first order incorporating non-zero contributions of the radion field and brane matter and this first order theory has a scalar-tensorial nature - this is also what we obtain by dimensional reduction [5] or by the technique of Shiromizu et al. [6].
@@
In this work, we will confine ourselves mostly to the two-brane Randall-Sundrum model which, along with its variants, have been thoroughly studied over the past two decades [7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24]. In its true form, the model has two opposite tension 3-branes embedded in a five-dimensional bulk spacetime and provides a possible resolve to the gauge hierarchy problem without introducing any intermediate scale in the theory. A central challenge in braneworld scenarios is the stabilization of the separation between the two branes—referred to as the modulus or radion. Achieving this requires generating an appropriate potential for the radion field that possesses a stable minimum, consistent with the value prescribed in the RS model, in order to resolve the gauge hierarchy problem. Goldberger and Wise (GW) [25] were the first to accomplish this by introducing a bulk scalar field endowed with suitable bulk and brane potentials. They demonstrated that the modulus can be stabilized without resorting to unnatural fine-tuning of parameters. Clearly, upon radion stabilization, the on-brane effective theory which was scalar-tensor (ST) earlier, becomes general relativity (GR) [26].
@@
Recently, the first-order thermodynamics of modified gravity theories has gained a lot of interest [27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45]. The primary idea is- if the effective fluid which appears in the modified Einstein equation of the corresponding gravity theory satisfies the Eckart-Fourier constitutive equation for dissipative fluids [46], then a “temperature” of the gravity theory can be defined consistently. Miraculously, one of such theories is scalar-tensor gravity, which has consequently led to a lot of exploration [27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42]. Another theory which passes the test is Horndeski gravity [43, 44, 45]. Much recently, Ref.[47] uses the first-order thermodynamic interpretation to formalize the attractor-to-GR mechanism in scalar-tensor theories. This was first proposed by Damour and Nordverdt [48, 49] and was subsequently plagued by a repellor mechanism [50] which only led to increasing complications in understanding the mechanism. Ref. [47] however provided a fresh perspective on the issue - it linked the relaxation of the effective fluid to its “zero temperature” equilibrium state to the attractor mechanism kicking in and taking scalar-tensor gravity towards GR, and vice versa. This formalism thus introduces the concept of a “gravitational temperature” along with an explicit equation governing the system’s evolution toward, or away from, GR equilibrium, making it particularly well-suited for studying the attractor-to-GR mechanism in cosmology.
@@
$S=\int\frac{d^{4}x}{16\pi}\sqrt{-g}\left[\phi R-\frac{\omega}{\phi}\,\nabla^{a}\phi\nabla_{a}\phi-V\right]+S^{\mathrm{(m)}}$ (1)
where $g_{ab}$ is the spacetime metric and $R$ is the Ricci scalar. The Brans-Dicke (BD) coupling $\omega>-3/2$ to avoid $\phi$ being a phantom field, and $S^{\mathrm{(m)}}$ is the matter action. The Einstein equations are then,
$R_{ab}-\frac{R}{2}\,g_{ab}=\frac{8\pi}{\phi}\,T_{ab}^{\mathrm{(m)}}+T_{ab}^{(\phi)}\,$ (2)
$\,\,\,T_{ab}^{(\phi)}=\frac{\omega}{\phi^{2}}\left(\nabla_{a}\phi\nabla_{b}\phi-\frac{1}{2}\,g_{ab}\nabla^{c}\phi\nabla_{c}\phi\right)$
$\qquad\quad\quad+\frac{1}{\phi}\left(\nabla_{a}\nabla_{b}\phi-g_{ab}\Box\phi\right)-\frac{V}{2\phi}g_{ab}\,\,$
$(2\omega+3)\Box\phi=8\pi T^{\mathrm{(m)}}+\phi\,V^{\prime}-2V-\omega^{\prime}\,\nabla^{c}\phi\nabla_{c}\phi\,$
where $R_{ab}$ is the Ricci tensor, $T_{ab}^{\mathrm{(m)}}$ is the matter energy-momentum tensor, $T^{(m)}=g^{ab}T_{ab}^{\mathrm{(m)}}$, $\nabla$ is the covariant derivative for Levi-Civita connection, $\Box=g^{ab}\nabla_{a}\nabla_{b}$, $T_{ab}^{(\phi)}$ is the effective energy-momentum tensor of $\phi$, and a prime denotes differentiation with respect to $\phi$.
With a timelike and future-oriented gradient $\nabla^{a}\phi$, the effective energy-momentum tensor of $\phi$ is equivalent to that of an effective dissipative fluid with 4-velocity $u^{a}:=\nabla^{a}\phi/\sqrt{-\nabla^{e}\phi\nabla_{e}\phi}$. The stress tensor of a dissipative imperfect fluid of 4-velocity $u^{a}$, in its full generality, can be written as
$T_{ab}=\rho u_{a}u_{b}+Ph_{ab}+\pi_{ab}+2\,q_{(a}u_{b)}\,$, (4)
where $\rho$, $P$, $\pi_{ab}$ and $q_{a}$ represent an effective energy density, effective isotropic pressure, effective anisotropic stress tensor and effective heat flux density respectively. $h_{ab}=g_{ab}+u_{a}u_{b}$ (describes hypersurface perpendicular to $u_{a}$), and the parentheses represents symmetrization. The fluid quantities associated with $T^{(\phi)}_{ab}$ are denoted by $\rho^{(\phi)}$, $P^{(\phi)}$, $\pi_{ab}^{(\phi)}$ and $q_{a}^{(\phi)}$. The most exciting observation that truly led to this alternate description of ST gravity is that this effective fluid obeys the Eckart-Fourier constitutive law for dissipative fluids [46] $q_{a}=-{\cal K}\left(\nabla_{a}{\cal T}+{\cal T}\dot{u}_{a}\right)$ where $\cal K$} is the thermal conductivity, $\cal T$} is the temperature, and $\dot{u}^{a}=u^{c}\nabla_{c}u^{a}$ is the fluid acceleration. A pedagogical calculation infact yields $q_{a}^{(\phi)}=-{\cal KT}\dot{u}_{a}$ where [47, 34, 35, 36]
@@
As one can clearly observe, ST gravity reduces to GR in the limit $\phi\to$ constant, in which ${\cal KT}\to 0$. The approach/departure of ST gravity to/from GR is described by [47, 34, 35, 36]
$\frac{d\left({\cal KT}\right)}{d\tau}=8\pi\left({\cal KT}\right)^{2}-\Theta{\cal KT}+\frac{\Box\phi}{8\pi\phi}\,$, (6)
where $\tau$ is the proper time along the effective fluid flow and $\Theta:=\nabla_{c}u^{c}$ is its expansion scalar. Substituting $\Box\phi$ from Eq.(3) gives us
@@
The induced metric on the positive (or negative) tension brane is thus $g_{(+)\mu\nu}=g_{\mu\nu}$ (or $g_{(-)\mu\nu}=e^{-2kr_{c}T}g_{\mu\nu}$), respectively. Since we impose ${\bf Z}_{2}$ symmetry ($z\leftrightarrow-z$), massless vector fluctuations associated with the off-diagonal part of the metric are absent [3, 51]. The low-energy effective theory can now be found by simple dimensional reduction (or by gradient expansion method as in Ref.[4] or by the technique of Ref.[6]).
@@
As we can see from Eq.(18), there are only two solutions with $\cal KT$} = constant ($\frac{d({\cal KT})}{d\tau}=0$). These are ${\cal KT}=0$ and ${\cal KT}=\Theta/8\pi(1+\frac{\omega_{\pm}}{3})$. Following the analysis of Ref.[47], we demonstrate this scenario on the $(\Theta,{\cal KT})$ plane (see Fig.1), albeit only the first quadrant. The quadrants where ${\cal KT}<0$ are discarded by definition. We also need not worry about the ($\Theta<0,{\cal KT}>0$) quadrant as Eq.(18) becomes: $\frac{d({\cal KT})}{d\tau}=8\pi(1+\frac{\omega_{\pm}}{3})({\cal KT})^{2}+|\Theta|{\cal KT}$. $(1+\frac{\omega_{\pm}}{3})$ being positive, this quadrant naturally takes away all solutions away from GR. The ${\cal KT}=\Theta/8\pi(1+\frac{\omega_{\pm}}{3})$ solution is identified as the critical curve - solutions lying below (above) it converge to (diverge from) GR. As the Brans-Dicke functions for both branes depend on the radion field, the critical curves are not straight lines. The comparative positions of critical curves on either brane can easily be seen from Eq.(19) - since $\frac{8\pi{\cal KT}}{\Theta}=\frac{1}{1+\frac{\omega_{+}}{3}}$, The Planck brane critical curve is below $8\pi{\cal KT}=\Theta$ line (this is representative of electrovacuum Brans-Dicke [47]) as $(1+\frac{\omega_{+}}{3})>1$ whereas the TeV brane critical curve is above the same as $\frac{1}{2}<(1+\frac{\omega_{-}}{3})<1$. Owing to the relative positions of the critical curves (see Fig.1), the TeV brane allows a larger class of solutions to flow towards GR than the Planck brane - in this sense, the TeV brane acts as a better “cooler” than the Planck brane.
@@
In case of an FLRW geometry on the branes with $a(t)$ being the scale factor, $\Theta=3H$ where $H=\frac{\dot{a}}{a}$. For the attractor mechanism to kick in, we need to have $8\pi{\cal KT}<\frac{\Theta}{1+\frac{\omega_{\pm}}{3}}$. Using the definition of $\cal KT$} from Eq.(5), we get $\frac{|\dot{\Phi}|}{\Phi}<\frac{3H}{1+\frac{\omega_{\pm}}{3}}$. In a contracting universe ($H<0$), for the attractor mechanism to work, $1+\frac{\omega_{\pm}}{3}$ has to be negative which is forbidden - hence, no attractor mechanism in such a universe. On the contrary, in an expanding universe ($H>0$), the attractor mechanism can indeed take place on both the branes provided that we satisfy
@@
Figure 1: A schematic representation of the critical curves in three specific theories - electrovacuum Brans-Dicke (blue) [see Ref.[47]], TeV brane Brans-Dicke (red) and Planck brane Brans-Dicke (green). For each of the theories, solutions starting above the critical curve deviate forever from GR, those starting below converge to GR. The critical curves cannot be crossed. On left - without a stabilizing field - dynamics about the stabilized radion. On right - with a stabilizing field.
@@
First let us break it down to two regimes: (a) $\Lambda_{4D}\rightarrow 0$: the 4D cosmological constant tuned to zero implies an almost flat Planck brane. This reduces the above combination to just the first term. Now, $\Phi_{+}<\alpha$ implies $(\frac{1}{\alpha}-\frac{1}{\Phi_{+}})<0$. Thus, the contribution now solely depends on the detuning of the visible brane tension. If $T_{vis}>\Lambda_{b}/k$, then the contribution is negative overall and it cools down gravity and vice versa. (b) $\tau\rightarrow 0$: This gets rid of the first term and the signature of the contribution depends only on the 4D cosmological constant i.e. detuning of the hidden brane tension. If $T_{hid}>-\Lambda_{b}/k$, then the contribution is negative implying a cooling down of gravity - and vice versa. The only thing we could decipher without pertaining to any regime is the following: if $T_{vis}>\Lambda_{b}/k$ and $T_{hid}>-\Lambda_{b}/k$ can be ensured simultaneously (which is certainly greater detuning), then the net contribution is negative - hence, making it easier for the theory to tend towards GR. But this in turn means that $T_{hid}+T_{vis}>0$. Recalling the Gibbons-Kallosh-Linde geometric consistency condition [55, 56],
@@
where $\oint$ is the integration over the internal compact space and $R_{p}$ is the Ricci scalar for the Planck brane metric. Thus, we can have the $T_{hid}+T_{vis}>0$ only if the Planck brane geometry is of the anti-de Sitter type ($R_{p}<0$).
@@
$\Phi_{-}>0$ and $\alpha>0$ - hence, their individual and additive contributions are all positive. We again consider the two regimes: (a) $\Lambda^{\prime}_{4D}\rightarrow 0\;(\tau\rightarrow 0)$: the 4D cosmological constant tuned to zero implies an almost flat TeV brane. This reduces the above combination to just the first term. Again, the contribution now solely depends on $\tau^{\prime}$ i.e. the detuning of the visible brane tension. If $T_{hid}>-\Lambda_{b}/k$, then the contribution is negative and it cools down gravity and vice versa. (b) $\tau^{\prime}\rightarrow 0\;(\Lambda_{4D}\rightarrow 0)$: This gets rid of the first term and the signature of the contribution depends only on the TeV brane 4D cosmological constant. If $T_{vis}>\Lambda_{b}/k$, then the contribution is negative cooling down gravity - and vice versa. Without restricting to a regime, the simultaneous detuning of both the brane tensions can ensure a negative contribution (as in the Planck brane case) but this is deemed possible only for an AdS Planck brane geometry as we argued previously.
A crucial observation is that in both the regimes (“almost flat Planck brane” and “almost flat TeV brane”), identical detuning of the brane tensions ensures identical signature for the radion potential contribution to the $\cal KT$} dynamics of the gravity theory (whether it “cools” or “heats” gravity) on either brane. For the $\Lambda_{4D}\rightarrow 0$ regime, the signature of $(T_{vis}-\Lambda_{b}/k)$ does the job; and for the $\Lambda^{\prime}_{4D}\rightarrow 0$ regime, the signature of $(T_{hid}+\Lambda_{b}/k)$ plays the equivalent role.
@@
Using $v_{h}/v_{v}=1.5$ and $m/k=0.2$ in Eq.(31) yields $kr_{c}T\approx 36$ and as one can see, no unnatural fine-tuning of parameters is required to solve the hierarchy problem [25]. Since the modulus gets stabilized to a fixed value, we can clearly see that the Brans-Dicke fields on both the branes which depend on the modulus field as
$e^{2kr_{c}T}=\frac{\alpha}{\alpha-\Phi_{+}}=1+\frac{\Phi_{-}}{\alpha}$ (32)
become constants. As a result, the theory on either brane eventually turns out to be GR (${\cal KT}=0$). The question is if we can witness the same from the thermal viewpoint i.e. the dynamical equation for $\cal KT$} without adhering to any particular geometry (like FLRW) on the brane. Accordingly, we plotted $V^{\prime}-\frac{2V}{\Phi_{\pm}}$ using Eqs. (30) and (32). The numerics were performed for the choice of parameters below Eq.(31) which do yield a value of the stabilized modulus deemed suitable for addressing the hierarchy problem. The result we get is the following: on both the branes, the signature of the $V^{\prime}-\frac{2V}{\Phi_{\pm}}$ contribution is negative till large values of $\Phi_{\pm}$ (on the Plack brane, the $\Phi_{+}$ axis is complete since $T\rightarrow\infty$ implies $\Phi_{+}\rightarrow\alpha$), hence cooling down the gravity theory on either brane (see Fig.2) - which is precisely what we expected to occur in the presence of a stabilizing field.
It is well known that a stabilizing field with negative mass term fails to stabilize the radion [54, 25] despite satisfying the Breitenlohner-Freedman bound.  [^1] In AdS space which is the nature of the bulk in Randall-Sundrum, the stability of the stabilizing field only requires satisfying the Breitenlohner-Freedman bound [57]: $m^{2}/k^{2}+4>0$ where $m^{2}$ is the mass squared of the GW field and $k$ is the AdS curvature scale. As the modulus never settles to a fixed value, we do not expect the gravity on either brane to flow towards GR. The same numerics were performed and the results show that the signature of $V^{\prime}-\frac{2V}{\Phi_{\pm}}$ on either brane eventually becomes positive for large enough values of $\Phi_{\pm}$. As the bulk field fails to stabilize the radion, the radion blows up (as the radion potential is ever decreasing) which indicates $\Phi_{+}$ reaching $\alpha$ and $\Phi_{-}\rightarrow\infty$. Hence, the presence of such a bulk field drives the increment of the radion field, which upon crossing a threshold renders the signature of $V^{\prime}-\frac{2V}{\Phi_{\pm}}$ positive (see Fig.3) - this drives the gravity theory on either brane away from GR.

### v2

@@ Abstract
A natural question to ask is- what is the low-energy effective theory on the branes? Since both cosmological and astrophysical solutions occur at energy scales much lower than that of the Planck scale, therefore in the effective theory approach the brane curvature radius $L$ is much large compared to bulk curvature $l$. As a result, perturbation theory can be used with a dimensionless perturbation parameter $\epsilon$ such that $\epsilon=(\frac{l}{L})^{2}<<1$. This method, called the Kanno-Soda gradient approximation scheme [50], is a metric-based iterative method in which the bulk metric and extrinsic curvature are expanded with increasing orders of $\epsilon$ in perturbation theory. The effective Einstein equations on a brane are determined with the solutions of these quantities and the junction conditions. When this method is applied to the two-brane Randall-Sundrum (RS) model [65], the RS fine-tuning condition (see Eq.9) is reproduced at the zeroth order when the radion is constant and the two $3$-branes are characterised by opposite brane tensions. The effective Einstein equations are then obtained at the first order incorporating non-zero contributions of the radion field and brane matter and this first order theory has a scalar-tensorial nature - this is also what we obtain by dimensional reduction [15].
In this work, we will confine ourselves mostly to the two-brane Randall-Sundrum model which, along with its variants, have been thoroughly studied over the past two decades [17, 55, 1, 24, 25, 23, 52, 62, 5, 7, 22, 3, 59, 56, 10, 19, 66, 68]. The two-brane RS scenario was originally introduced as a geometric resolution of the gauge hierarchy problem, providing an exponential suppression of mass scales on the visible brane without introducing an additional large fundamental hierarchy. This was in stark contrast to large extra-dimensional models, where the hierarchy reappears in the form of a very large compactification radius. Although direct experimental signatures such as Kaluza-Klein (KK) gravitons in collider searches or radion-induced cosmological effects have not yet been detected, it remains plausible that they may just be eluding the present energy scale, since even a $\mathcal{O}(100TeV)$ KK graviton, which will be beyond the present energy scale, may result from a very little hierarchy introduced between the bulk AdS scale and the radius of the orbifolded compact dimension. From a theoretical standpoint, the RS warped background closely parallels the “throat” geometries that arise in string theory motivated compactifications. Also as we discussed, RS is one of the fundamental models yielding a scalar-tensor effective theory whose necessity in our further analysis is indispensable.
@@
In its true form, the model has two opposite tension 3-branes embedded in a five-dimensional bulk spacetime. A central challenge in braneworld scenarios is the stabilization of the separation between the two branes - referred to as the modulus or radion. Achieving this requires generating an appropriate potential for the radion field that possesses a stable minimum, consistent with the value prescribed in the RS model, in order to resolve the gauge hierarchy problem. Goldberger and Wise (GW) [47] were the first to accomplish this by introducing a bulk scalar field endowed with suitable bulk and brane potentials. They demonstrated that the modulus can be stabilized without resorting to unnatural fine-tuning of parameters. Clearly, upon radion stabilization, the on-brane effective theory which was scalar-tensor (ST) earlier, becomes general relativity (GR) [18].
@@
Recently, the first-order thermodynamics of modified gravity theories has gained a lot of interest [42, 38, 39, 31, 30, 45, 40, 41, 33, 32, 51, 49, 29, 35, 58, 28, 57, 46, 36]. The primary idea is- if the effective fluid which appears in the modified Einstein equation of the corresponding gravity theory satisfies the Eckart-Fourier constitutive equation for dissipative fluids [26], then a “temperature” of the gravity theory can be defined consistently. Miraculously, one of such theories is scalar-tensor gravity, which has consequently led to a lot of exploration [42, 38, 39, 31, 30, 45, 40, 41, 33, 32, 51, 49, 29, 35, 58, 28]. More generally, Horndeski theories pass the same test [57, 46, 36]. Much recently, Ref.[34] uses the first-order thermodynamic interpretation to formalize the attractor-to-GR mechanism in scalar-tensor theories. This was first proposed by Damour and Nordverdt [20, 21] and was subsequently plagued by a repellor mechanism [67] which only led to increasing complications in understanding the mechanism. Ref. [34] however provided a fresh perspective on the issue - it linked the relaxation of the effective fluid to its “zero temperature” equilibrium state to the attractor mechanism kicking in and taking scalar-tensor gravity towards GR, and vice versa. This formalism thus introduces the concept of a “gravitational temperature” along with an explicit equation governing the system’s evolution toward, or away from, GR equilibrium, making it particularly well-suited for studying the attractor-to-GR mechanism in cosmology.
@@
$S=\int\frac{d^{4}x}{16\pi}\sqrt{-g}\left[\phi R-\frac{\omega(\phi)}{\phi}\,\nabla^{a}\phi\nabla_{a}\phi-V\right]+S_{m}$ (1)
with $g_{ab}$ being the metric and $R$ the Ricci scalar. The Brans-Dicke (BD) function $\omega>-3/2$ to avoid $\phi$ being a phantom field, and $S_{m}$ is the matter action.
With a timelike and future-oriented gradient $\nabla^{a}\phi$, the effective energy-momentum tensor $T_{ab}^{(\phi)}$ of $\phi$ is equivalent to that of an effective dissipative fluid with 4-velocity $u^{a}:=\nabla^{a}\phi/\sqrt{-\nabla^{e}\phi\nabla_{e}\phi}$. The stress tensor of a dissipative imperfect fluid of 4-velocity $u^{a}$, in its full generality, can be written as
$T_{ab}=\rho u_{a}u_{b}+Ph_{ab}+\pi_{ab}+q_{a}u_{b}+q_{b}u_{a}$ (2)
where $\rho$, $P$, $\pi_{ab}$ and $q_{a}$ represent an effective energy density, effective isotropic pressure, effective anisotropic stress tensor and effective heat flux density respectively. $h_{ab}=g_{ab}+u_{a}u_{b}$ (describes hypersurface perpendicular to $u_{a}$). The fluid quantities associated with $T^{(\phi)}_{ab}$ are denoted with a $(\phi)$ superscript. The observation that truly led to this alternate description of ST gravity is that this effective fluid obeys the Eckart-Fourier constitutive law for dissipative fluids [26] $q_{a}=-{\cal K}\left(\nabla_{a}{\cal T}+{\cal T}\dot{u}_{a}\right)$ where $\cal K$} is the thermal conductivity, $\cal T$} is the temperature, and $\dot{u}^{a}=u^{c}\nabla_{c}u^{a}$ is the fluid acceleration. A pedagogical calculation infact yields $q_{a}^{(\phi)}=-{\cal KT}\dot{u}_{a}$ where [34, 41, 33, 32]
@@
As we observe, ST gravity reduces to GR in the limit $\phi\to$ constant i.e. ${\cal KT}\to 0$. The convergence (divergence) of ST gravity to (from) GR is described by [34, 41, 33, 32]
$\frac{d\left({\cal KT}\right)}{d\tau}=8\pi\left({\cal KT}\right)^{2}-\Theta{\cal KT}+\frac{\Box\phi}{8\pi\phi}\,$ (4)
where $\tau$ is the proper time along the $\phi$ fluid flow and $\Theta:=\nabla_{c}u^{c}$ is its expansion scalar. Incorporating the scalar equation of motion to substitute for $\square\phi$, we get
@@
The induced metric on the positive (or negative) tension brane is thus $g_{(+)\mu\nu}=g_{\mu\nu}$ (or $g_{(-)\mu\nu}=e^{-2kr_{c}T}g_{\mu\nu}$), respectively. Since we impose ${\bf Z}_{2}$ symmetry ($z\leftrightarrow-z$), massless vector fluctuations associated with the off-diagonal part of the metric are absent [65, 48]. The low-energy effective theory can now be found by simple dimensional reduction (or by gradient expansion method as in Ref.[50]).
@@
As we can see from Eq.(16), there are only two solutions with $\cal KT$} = constant ($\frac{d({\cal KT})}{d\tau}=0$). These are ${\cal KT}=0$ and ${\cal KT}=\Theta/8\pi(1+\frac{\omega_{\pm}}{3})$. Following the analysis of Ref.[34], we demonstrate this scenario on the $(\Theta,{\cal KT})$ plane (see Fig.1), albeit only the first quadrant. The quadrants where ${\cal KT}<0$ are discarded by definition. We also need not worry about the ($\Theta<0,{\cal KT}>0$) quadrant as Eq.(16) becomes: $\frac{d({\cal KT})}{d\tau}=8\pi(1+\frac{\omega_{\pm}}{3})({\cal KT})^{2}+|\Theta|{\cal KT}$. $(1+\frac{\omega_{\pm}}{3})$ being positive, this quadrant naturally takes away all solutions away from GR. The ${\cal KT}=\Theta/8\pi(1+\frac{\omega_{\pm}}{3})$ solution is identified as the critical curve - solutions lying below (above) it converge to (diverge from) GR. As the Brans-Dicke functions for both branes depend on the radion field, the critical curves are not straight lines. The comparative positions of critical curves on either brane can easily be seen from Eq.(17) - since $\frac{8\pi{\cal KT}}{\Theta}=\frac{1}{1+\frac{\omega_{+}}{3}}$, the Planck brane critical curve is below $8\pi{\cal KT}=\Theta$ line (this is representative of electrovacuum Brans-Dicke [34]) as $(1+\frac{\omega_{+}}{3})>1$ whereas the TeV brane critical curve is above the same as $\frac{1}{2}<(1+\frac{\omega_{-}}{3})<1$. Owing to the relative positions of the critical curves (see Fig.1), the TeV brane allows a larger class of solutions to flow towards GR than the Planck brane - in this sense, the TeV brane acts as a better “cooler” for ST theories than the Planck brane. We should remember that owing to the dependence of both the BD scalars on the radion field (see Eq.12), the BD scalar on one brane going to a constant value automatically implies the same for the other brane’s BD scalar - hence, the signature of $\frac{d(\cal KT)}{d\tau}$ is identical on either brane. Thus, the apparent better cooling nature of the TeV brane stems from the parametric dependence of the BD scalars on the radion field as
$e^{2kr_{c}T}=\frac{\alpha}{\alpha-\Phi_{+}}=1+\frac{\Phi_{-}}{\alpha}$ (18)
As one can observe, $\Phi_{-}$ is unbounded above whereas $\Phi_{+}$ can only go up to $\alpha$, thereby explaining the disparity.
@@
In case of an FLRW geometry on the branes with $a(t)$ being the scale factor, $\Theta=3H$ where $H=\frac{\dot{a}}{a}$. For the attractor mechanism to kick in, we need to have $8\pi{\cal KT}<\frac{\Theta}{1+\frac{\omega_{\pm}}{3}}$. Using the definition of $\cal KT$} from Eq.(3), we get $\frac{|\dot{\Phi}|}{\Phi}<\frac{3H}{1+\frac{\omega_{\pm}}{3}}$. In a contracting universe ($H<0$), for the attractor mechanism to work, $(1+\frac{\omega_{\pm}}{3})$ has to be negative which is forbidden - hence, no attractor mechanism in such a universe. On the contrary, in an expanding universe ($H>0$), the attractor mechanism can indeed take place on both the branes provided that we satisfy
@@
Braneworld models are commonly associated with ekpyrotic and cyclic cosmology [54]. The model postulates an attractive force between the two branes that causes them to slowly approach one another, leading to an ekpyrotic phase in which the universe undergoes a slow contracting stage that flattens the branes to a high degree aided by the on-brane ST theory. Eventually, the branes collide, and from the perspective of observers on a brane this collision appears as the Big Bang. After the collision, the brane separation nearly stabilizes, but a small residual attraction remains, acting like quintessence and providing the observed dark energy, and over very long timescales the attraction strengthens again, leading to another ekpyrotic phase, another collision, and thus a cyclic cosmology with repeated periods of contraction, collision, and expansion. In such cosmologies, the thermal description for an FLRW geometry tells us that for the contracting phase ($H<0$), no attractor-to-GR mechanism is possible but it may again get activated in the subsequent expanding phase ($H>0$). Similar analysis can be carried out for more general geometries.
Figure 1: A schematic representation of the critical curves in three specific theories - electrovacuum Brans-Dicke (blue) [see Ref.[34]], TeV brane Brans-Dicke (red) and Planck brane Brans-Dicke (green). For each of the theories, solutions starting above the critical curve deviate forever from GR, those starting below converge to GR. The critical curves cannot be crossed. On left - without a stabilizing field. On right - with a stabilizing field - dynamics about the stabilized radion.
@@
Let us break it down to two regimes: (a) $\Lambda_{4D}\rightarrow 0$: the 4D cosmological constant tuned to zero implies an almost flat Planck brane. This reduces the above combination to just the first term. $\Phi_{+}<\alpha$ implies $(\frac{1}{\alpha}-\frac{1}{\Phi_{+}})<0$. Thus, the contribution now solely depends on the detuning of the visible brane tension. If $T_{vis}>\Lambda_{b}/k$, then the contribution is negative overall and it cools down gravity and vice versa. (b) $\tau\rightarrow 0$: This gets rid of the first term and the signature of the contribution depends only on the 4D cosmological constant i.e. detuning of the hidden brane tension. If $T_{hid}>-\Lambda_{b}/k$, then the contribution is negative implying a cooling down of gravity - and vice versa. The only thing we could decipher without pertaining to any regime is the following: if $T_{vis}>\Lambda_{b}/k$ and $T_{hid}>-\Lambda_{b}/k$ can be ensured simultaneously (which is certainly greater detuning), then the net contribution is negative - hence, making it easier for the theory to tend towards GR. But this in turn means that $T_{hid}+T_{vis}>0$. Recalling the Gibbons-Kallosh-Linde geometric consistency condition [43, 53],
@@
where $\oint$ is the integration over the internal compact space and $R_{p}$ is the Ricci scalar for the Planck brane metric. Thus, we can have the $T_{hid}+T_{vis}>0$ only if the Planck brane geometry is of the anti-de Sitter type ($R_{p}<0$). Similarly, heating the theory would require a de-Sitter Planck brane geometry ($R_{p}>0$).
@@
One can observe Eq.(25) to reach the same conclusions as in the Planck brane case i.e. in the different regimes, the same detunings yield the same signatures for the contribution  [^1] The physical meaning of $\tau\rightarrow 0$ is $\Lambda^{\prime}_{4D}\rightarrow 0$ i.e. an almost flat TeV brane.. This is in fact an outcome of the stronger result we obtain upon conformal tansformation,
$\frac{dV}{d\Phi_{+}}-\frac{2V}{\Phi_{+}}=\frac{dV}{d\Phi_{-}}-\frac{2V}{\Phi_{-}}$ (26)
A crucial observation is that in both the regimes (“almost flat Planck brane” and “almost flat TeV brane”), non-dynamical detunings of the brane tensions entirely determine the signature for the radion potential contribution to the $\cal KT$} dynamics of the gravity theory (whether it “cools” or “heats” gravity) on either brane. For the $\Lambda_{4D}\rightarrow 0$ regime, the signature of $(T_{vis}-\Lambda_{b}/k)$ does the job; and for the $\Lambda^{\prime}_{4D}\rightarrow 0$ regime, the signature of $(T_{hid}+\Lambda_{b}/k)$ plays the equivalent role.
@@
One should note that in the absence of a stabilization mechanism, the radion remains massless but not tachyonic, reflecting a flat direction rather than an instability of the background [65, 15, 47, 48]. The transverse traceless graviton perturbations reduce to a Schrodinger type eigenvalue problem with the usual “volcano” potential, yielding a normalizable zero-mode reproducing four-dimensional gravity together with a KK tower satisfying $m^{2}>0$ [65, 64, 19, 66]. Thus, the absence of a stabilizing field does not by itself generate geometric pathologies or tachyonic KK instabilities - instead, the inter-brane separation remains dynamically undetermined.
@@
Using $v_{h}/v_{v}=1.5$ and $m/k=0.2$ in Eq.(31) yields $kr_{c}T\approx 36$ and as one can see, no unnatural fine-tuning of parameters is required to solve the hierarchy problem [47]. Since the modulus gets stabilized to a fixed value, we can clearly see that the Brans-Dicke fields on both the branes which depend on the modulus field (Eq.12) become constants. As a result, the theory on either brane eventually turns out to be GR (${\cal KT}=0$). The question is if we can witness the same from the thermal viewpoint i.e. the dynamical equation for $\cal KT$} without adhering to any particular geometry (like FLRW) on the brane. Accordingly, we plotted $V^{\prime}-\frac{2V}{\Phi_{\pm}}$ using Eqs. (12) and (30). The numerics were performed for the choice of parameters below Eq.(31) which do yield a value of the stabilized modulus deemed suitable for addressing the hierarchy problem. The result we get is the following: on both the branes, the signature of the $V^{\prime}-\frac{2V}{\Phi_{\pm}}$ contribution is negative till large values of $\Phi_{\pm}$ (on the Plack brane, the $\Phi_{+}$ axis is complete since $T\rightarrow\infty$ implies $\Phi_{+}\rightarrow\alpha$), hence cooling down the gravity theory on either brane (see Fig.2) - which is precisely what we expect to occur in the presence of a stabilizing field.
It is well known that a stabilizing field with negative mass term fails to stabilize the radion [8, 47] despite satisfying the Breitenlohner-Freedman bound.  [^2] In AdS space which is the nature of the bulk in Randall-Sundrum, the stability of the stabilizing field only requires satisfying the Breitenlohner-Freedman bound [9]: $m^{2}/k^{2}+4>0$ where $m^{2}$ is the mass squared of the GW field and $k$ is the AdS curvature scale. As the modulus never settles to a fixed value, we do not expect the gravity on either brane to flow towards GR. The same numerics were performed and the results show that the signature of $V^{\prime}-\frac{2V}{\Phi_{\pm}}$ on either brane eventually becomes positive for large enough values of $\Phi_{\pm}$. As the bulk field fails to stabilize the radion, the radion blows up (as the radion potential is ever decreasing) which indicates $\Phi_{+}$ reaching $\alpha$ and $\Phi_{-}\rightarrow\infty$. Hence, the presence of such a bulk field drives the increment of the radion field, which upon crossing a threshold renders the signature of $V^{\prime}-\frac{2V}{\Phi_{\pm}}$ positive (see Fig.3) - this drives the gravity theory on either brane away from GR  [^3] One may worry that the TeV brane scalar is unbounded - hence commenting on the contribution’s signature throughout is unjustified. However, this is not the case as Eq.(26) tells us that the contribution is identical on both branes and the entire range of the Planck brane BD scalar has been explored, providing a full knowledge of the TeV brane contribution.